Crew 188 Sol Summary Report 30JAN2018
Summary Title: The Human Factor
Author’s name: Renee Garifi, MDRS Crew 188 Executive Officer Mission Status: On track
Sol Activity Summary:
We awoke this morning to a beautiful Martian sunrise and amazing view of the alien landscape out the port hole windows on the second floor of the hab. We held our daily planning conference over breakfast where we mapped out our plan for the day including EVA prep, research hardware set up and how best to accomplish our daily tasks. We also spent some time checking in with each crewmember to discuss factors affecting our health, stress and personal mission goals. We had a positive discussion and everyone felt heard, which is important to the leadership of this team.
We began EVA prep at 09:30 and sent 3 members of our team through the airlock to explore the Martian terrain while the 3 remaining crew set to work setting up research hardware in the greenhouse and science lab. The Artist in Residence and Crew Astronomer have been working all day to prepare a special 360 degree video of the Super Blue Blood Moon Eclipse which begins tonight. Our best viewing opportunity will begin at 04:44 am.
One very exciting event occurred after our EVA crew returned. We received a cargo shipment of our embroidered mission patches! This provided our crew with an unexpected morale boost and helped us feel that much more excited about carrying out a high fidelity analog mission worthy of an official crew mission patch.
MDRS Crew 188 Executive Officer
Look Ahead Plan:
In the late evening/early morning tomorrow, all crewmembers will rise at 04:00 am to view the special astronomical mega event of the lunar eclipse of the January Blue Moon.
Anomalies in work:
MDRS Director is coordinating power system updates with contractors. Robotic observatory currently not functional. MDRS Astronomy lead working problem.
A warm day on “Mars” with crewmembers getting quite a lot of sun on today’s EVA.
Crew Physical Status:
Crew are fighting fatigue and dehydration symptoms but are in good spirits.
One EVA successfully completed today. We will not have an EVA tomorrow (31 January).
Reports to file:
1. Sol Summary
2. Operations Report
3. Journalist Report
4. EVA #3 Report
6. Daily Photos
7. Mission Summary
8. Green Hab Report
9. Creative Report
Support Requested: See Operations Report